OpenStudy (richgall____):

I will medal and fan or whatever i just need this one question done today....

OpenStudy (richgall____):

A cylinder and a cone have the same diameter: 8 inches. The height of the cylinder is 5 inches. The height of the cone is 15 inches. Use pi = 3.14. What is the relationship between the volume of this cylinder and this cone? Explain your answer by determining the volume of each and comparing them. Show all your work.

OpenStudy (richgall____):

@daniel12900 @AlexandervonHumboldt2 @Tamara7301 @mjmahmood @green_1 @welshfella

OpenStudy (tamara7301):

@imqwerty

OpenStudy (green_1):

i think it is 301.44 but im not sure. @mathmate

imqwerty (imqwerty):

 1st of all diameter=8 so radius=4 cuz radius is half of diameter \[volume_{cylinder}=pi \times r^2 \times h\]\[volume_{cylinder}=3.14\times 4^2 \times 5\] \[volume_{cone}=\frac{ \pi \times r^2 \times h }{ 3 }\]\[volume_{cone}=\frac{ 3.14 \times 4^2 \times 15 }{ 3 }\]\[volume_{cone}=3.14 \times 4^2 \times 5\] clearly volume of cylinder=volume of cone